# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v2

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adaptations (LoRAs) has become common practice, often yielding numerous copies of the same LLM differing only in their LoRA updates. This paradigm presents challenges for systems that serve real-time responses to queries that each involve a different LoRA. Prior works optimize the design of such systems but still require continuous loading and offloading of LoRAs, as it is infeasible to store thousands of LoRAs in GPU memory. To mitigate this issue, we investigate the efficacy of model compression when serving LoRAs. We propose a method for joint compression of LoRAs into a shared basis paired with LoRA-specific scaling matrices. We extend our algorithm to learn clusters of LoRAs that are more amenable to joint compression, allowing it to scale gracefully to large LoRA collections. Our experiments with up to 500 LoRAs demonstrate that compressed LoRAs preserve performance while offering major throughput gains in realistic serving scenarios with over a thousand LoRAs, maintaining 80% of the throughput of serving a single LoRA.
1MIT CSAIL, 2MIT-IBM Watson AI Lab
@@ 1 Introduction
The myriad uses for foundation models (FMs) have led to a proliferation of specialized models, each fine-tuned to perform a downstream task. To avoid fine-tuning foundation models with billions of parameters, more parameter-efficient fine-tuning (PEFT) algorithms were proposed. An especially successful PEFT method is low-rank adaptation (LoRA) (Hu et al., 2021), which learns low-rank additive changes to neural network matrices. Because of the low-rank parameterization, these matrices (called adapter weights) contain orders-of-magnitude fewer parameters than the base model. Still, LoRA can achieve performance on par with full fine-tuning (Hu et al., 2021).
LoRA’s popularity has triggered a growing need to serve large collections of LoRA adapters at scale. For example, proprietary and open-source LLM providers offer fine-tuning services (OpenAI, 2024; TogetherAI, 2024; Predibase, 2024) with user bases likely in thousands or even hundreds of thousands. As each user wants to use their own fine-tuned version of the LLM, simply serving a dedicated fine-tuned LLM per user becomes infeasible. To this end, S-LoRA (Sheng et al., 2023) considers a system where only the base LLM is placed on an inference server and individual LoRA adapters are switched as needed at inference time. S-LoRA optimizes the system’s inner workings via custom CUDA kernels and memory management to increase the throughput when serving multiple LoRAs. Such multi-LoRA system design has also been adopted in vLLM (Kwon et al., 2023), a state-of-the-art LLM serving engine. Despite the optimized system design, serving LoRAs still has a fundamental limitation: when the number of adapters is large, they need to be constantly loaded and offloaded from GPU memory to accommodate incoming requests, degrading throughput.
The problem of accommodating multiple LoRA adapters is also apparent when placing LLMs on edge devices, where smaller LLMs are fine-tuned for various tasks and the adapters are swapped depending on the task at hand (Gunter et al., 2024). In this case, the amount of adapters is smaller, e.g., a few dozen (Gunter et al., 2024), but the memory constraints are also tighter due to the limited capacity of edge devices.
In this work, we consider the problem of compressing a collection of LoRAs. We have two key objectives: (1) preserving the performance of the original LoRAs and (2) improving the throughput of serving many LoRAs. We formulate LoRA compression as a reconstruction problem, where the goal is to approximate the original adapters via collections of matrices of a smaller total size. We investigate an approach based on compressing LoRAs jointly by finding a shared basis and LoRA-specific scaling matrices, and propose a joint diagonalization-based algorithm (JD). To improve reconstruction error for large numbers of LoRAs while keeping the number of parameters in check, we propose a clustering approach where each cluster is compressed independently using the joint diagonalization algorithm. Our clustering algorithm is based on alternating between optimizing the cluster assignments and the per-cluster reconstruction error.
@@
We showcase the benefits of joint compression in Figure 1. We chose JD configurations that match the constraints. When serving up to 64 unique LoRAs we use JD without clustering and for 128 or more, we pick the number of clusters to match the performance of compressed and original LoRAs. In each case, the GPU memory footprint of the compressed and original LoRAs is matched for a fair comparison to the vLLM’s multi-LoRA inference engine. When serving over 1000 LoRAs, compression increases throughput by a factor of 1.6 and maintains 80% of the throughput of serving the base LLM (or a single LoRA merged into the LLM). Detailed results are presented in Section 6.
@@
  • For large numbers of LoRAs, we scale the joint compression scheme by proposing a clustering algorithm where each cluster is jointly compressed to minimize the overall reconstruction error.
  • We establish theoretical guarantees for the reconstruction error central to our compression formulation and verify the relation between reconstruction loss and performance empirically.
  • We train a collection of 500 high-quality LoRAs for Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) on 500 natural instruction tasks (Wang et al., 2022) and demonstrate that our compression techniques preserve the performance of the original LoRAs. We will release the 500 LoRAs to facilitate future work on LoRA compression as well as the code for our method.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT) has become prevalent for updating foundation models thanks to the need for efficiency in training and communication (Lialin et al., 2023). Many PEFT methods have been proposed, e.g. (Houlsby et al., 2019; Liu et al., 2022) and LoRA (Hu et al., 2021) became the standard, partially due to the ease of switching between LoRAs in inference time.
Several works improve LoRA (Liu et al., 2024; Wang et al., 2024), sometimes with algebraic methods like SVD (Meng et al., 2024; Zhang et al., 2023; Jiang et al., 2023b) or by leveraging its statistical properties (Zhu et al., 2024; Zeng & Lee, 2024). Relatively few, however, accelerate inference times. S-LoRA (Sheng et al., 2023) provides an efficient means of switching between LoRAs. F-LoRA (Wen & Chaudhuri, 2024) adapts training to reduce batch multiplications, accelerating inference. Our method achieves a similar outcome (see Appendix D) without changing the LoRA formulation or requiring that LoRAs be trained in a dedicated way; future improvements to LoRA will also benefit from this aspect of our work (e.g., Meng et al. (2024)).
There are many efforts to compress models (Cheng et al., 2017; Gholami et al., 2022; Sharma et al., 2024; Li et al., 2018)—including some specifically for LoRAs—to accelerate inference. Predominantly, pruning and sparsification methods delete weights (Yadav et al., 2023a), and quantization methods reduce the weights’ precision (Dettmers et al., 2024). Some works compress weights to reduce model size but typically require decompression and hence do not save GPU memory (Hershcovitch et al., 2024). Similarly to our work, while most methods increase speed at the cost of performance, a few note increased performance and generalization after compression (Yadav et al., 2023a; Nadjahi et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2024).
Our work also relates to model merging (Choshen et al., 2022; Wortsman et al., 2022; Matena & Raffel, 2021) and mixture of experts methods (Muqeeth et al., 2024; Yadav et al., 2024). These methods reuse models trained by others (Choshen et al., 2023; Raffel, 2023), serving them together as one compressed model. Despite this similarity, these methods create a single general model that acts on any input, while our model allows for more performant per-task solutions.
@@ 3 Rank-Based LoRA Compression
While our compression technique has access only to a collection of $\{(A_{i},B_{i})\}_{i=1}^{n}$ pairs, in our experiments we will assess the efficacy of compression by comparing how the compressed matrices perform relative to the uncompressed LoRAs on typical data. For this reason, although in this section we optimize a Frobenius norm reconstruction error relative to the product $B_{i}A_{i}$, in reality this is a proxy for the nonlinear and complex way that compression errors in the adapters impact transformer performance. Our experimental evaluation will thus focus on the performance of the compressed LoRAs against the uncompressed versions on real data in §6.
Our compression methods significantly reduce the overall number of parameters. Reducing parameters through compression theoretically accelerates storage and serving processes for a collection of LoRAs. This reduction, however, alters the computational dynamics during inference, so parameter reduction alone does not immediately imply faster throughput. In light of the complexities of GPU optimization, we experimentally assess the throughput under realistic conditions in §6.3.
@@ 3.1 Joint Diagonalization
For compression to scale to large numbers of LoRAs, the compressed number of parameters should not scale linearly with $n$. Hence seeking to compress each LoRA individually (e.g., via SVD as detailed in the experimental baselines) is inherently limited.
To address this, we suggest a Joint Diagonalization (JD) method, which optimizes a shared basis onto which we can project the set of $n$ LoRAs. This will allow structure to be shared, implicitly grouping and/or merging the collection of LoRAs.
In this model, each LoRA product $B_{i}A_{i}$ is factorized into the form $U\Sigma_{i}V$, where $U$ and $V$ are shared across all LoRAs and $\Sigma_{i}$ is specific to each LoRA. In this formulation, every $\Sigma_{i}$ shares the same rank $r$. This allows $U$ and $V$ to be pre-loaded onto the GPU, with $\Sigma_{i}$ loaded when necessary for each batch. The matrices $\Sigma_{i}$ can be either diagonal or small square matrices, accelerating the forward pass compared to conventional multi-LoRA serving configurations.
Objective function. Motivated by the relationship of singular value decomposition to minimizing the Frobenius norm of the reconstruction error, we also propose to minimize the Frobenius norm of the adapter matrix approximation error. Specifically, we use the following objective function:
@@
Full $\Sigma_{i}$ approximation. The first method we call $\mathsf{JD-Full}$. Without loss of generality, $U$ and $V$ can be constrained to be orthogonal, so long as $\Sigma_{i}$ remains an unconstrained full matrix. $\mathsf{JD-Full}$ adopts this restriction to make the optimization better posed, but note it does not restrict the expressiveness of the objective equation 1. This setting yields the following optimization problem:
$\operatorname{JD-Full}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})=\underset{U^{\top}U=V^{\top}V=I_{r}}{\underset{\{\Sigma_{i}\}_{i=1}^{n}}{\operatorname{argmin}}}\sum_{i=1}^{n}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}~{}~{}~{}~{}(\mathsf{JD-Full})$ (2)
An efficient alternating algorithm to solve this objective function can be found in Appendix A.
@@
$\operatorname{JD-Diag}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})={\underset{\{\Sigma_{i}\}_{i=1}^{n},U,V}{\operatorname{argmin}}}\sum_{i=1}^{n}\|B_{i}A_{i}-U\mathrm{diag}(\Sigma_{i})V^{\top}\|_{\mathrm{Fro}}^{2}~{}~{}~{}~{}(\mathsf{JD-Diag})$ (3)
An efficient alternating least squares algorithm to optimize this objective can be found in Appendix A. This diagonal version has some per-LoRA parameter savings when compared to $\mathsf{JD-Full}$, since the diagonal $\Sigma_{i}$ only needs $r$ parameters instead of $r^{2}$.
@@ 3.2 Clustering
As the number of LoRAs $n$ grows and becomes more diverse, the rank $r$ needed for Joint Diagonalization to achieve good performance will tend to increase. This increases the size and number of parameters of each $\Sigma_{i}$ that needs to be stored, especially for $\operatorname{JD-Full}$ which will require $O(nr^{2})$ storage for these matrices. If the necessary $r$ is growing proportionally to $n$, then this storage will eventually become the bottleneck.
To resolve this limitation with very large $n$, we propose to group the $n$ LoRAs into $|C_{j}|$ clusters $C_{j}$. Each cluster is given its own rank $r$ JD compression, and the clusters are chosen such that the overall reconstruction error is minimized. Specifically, the overall objective is
$\min_{\{\{C_{j}\},{U_{j},V_{j}}\},\{\Sigma_{i}\}}\sum_{j}\sum_{i\in C_{j}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}||^{2}_{F},$
and we optimize this by alternating between cluster assignments and the JD of each cluster. The algorithm details are in Appendix A.3. Typically, the goal with large $n$ is to have $|C_{j}|$ grow with $n$ as $r$ becomes fixed. Comparing $k$ rank-$r$ $\operatorname{JD-Full}$ clusters to a rank-$kr$ $\operatorname{JD-Full}$ single cluster compression, observe that the clustered approach requires $O(dkr+nr^{2})$ parameters, while the single-cluster approach requires $O(dkr+nk^{2}r^{2})$ parameters due to the increased sizes of the $\Sigma_{i}$s. While these two approaches have the same rank, note that they may have different reconstruction abilities. Empirically, we find that multiple clusters significantly aid performance for $n\geq 100$.

### v3

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adaptations (LoRAs) has become common practice, often yielding numerous copies of the same LLM differing only in their LoRA updates. This paradigm presents challenges for systems that serve real-time responses to queries that each involve a different LoRA. Prior works optimize the design of such systems but still require continuous loading and offloading of LoRAs, as it is infeasible to store thousands of LoRAs in GPU memory. To mitigate this issue, we investigate the efficacy of compression when serving LoRAs. We propose a method for the joint compression of LoRAs into a shared basis paired with LoRA-specific scaling matrices. We extend our algorithm to learn clusters of LoRAs that are amenable to joint compression, allowing it to scale gracefully to large LoRA collections. Our experiments with up to 1000 LoRAs [^1] https://huggingface.co/Lots-of-LoRAs demonstrate that compressed LoRAs preserve performance while offering major throughput gains in realistic serving scenarios with over a thousand LoRAs, maintaining 80% of the throughput of serving a single LoRA.
Machine Learning, ICML
@@ 1 Introduction
The myriad uses for foundation models (FMs) have led to a proliferation of specialized models, each fine-tuned to perform a downstream task. To avoid fine-tuning foundation models with billions of parameters, parameter-efficient fine-tuning (PEFT) algorithms were proposed. An especially successful PEFT method is low-rank adaptation (LoRA) (Hu et al., 2021), which learns low-rank additive changes to neural network matrices. Because of the low-rank parameterization, these matrices (called adapter weights) contain orders of magnitude fewer parameters than the base model. Still, LoRA can achieve performance on par with full fine-tuning (Hu et al., 2021).
@@
LoRA’s popularity has triggered a growing need to serve large collections of LoRA adapters at scale. Proprietary and open-source LLM providers offer fine-tuning services (OpenAI, 2024; TogetherAI, 2024; Predibase, 2024) with user bases likely in the thousands or even hundreds of thousands. As each user wants to use their own fine-tuned version of the LLM, serving a dedicated fine-tuned LLM per user becomes infeasible. To this end, S-LoRA (Sheng et al., 2023) proposes a system where only the base LLM is placed on an inference server and individual LoRA adapters are switched as needed at inference time. S-LoRA optimizes the system’s inner workings via custom CUDA kernels and memory management to increase throughput when serving multiple LoRAs. Multi-LoRA system design has also been adopted in vLLM (Kwon et al., 2023), a state-of-the-art LLM serving engine. Despite optimized system designs, serving LoRAs still has a fundamental limitation: when the number of adapters is large, they need to be constantly loaded and offloaded from GPU memory to accommodate incoming requests, degrading throughput.
The problem of accommodating multiple LoRA adapters is also apparent when placing LLMs on edge devices, where smaller LLMs are fine-tuned for various tasks, and the adapters are swapped depending on the task at hand (Gunter et al., 2024). In this setting, the number of adapters is smaller, e.g., a few dozen (Gunter et al., 2024), but the memory constraints are also more stringent.
In this work, we consider the problem of compressing a collection of LoRAs. We have two key objectives: (1) preserving the performance of the original LoRAs and (2) improving the throughput of serving many LoRAs. We formulate LoRA compression as a reconstruction problem, where the goal is to approximate the original adapters via collections of matrices of a smaller size. We compress LoRAs jointly by finding a shared basis and LoRA-specific scaling matrices and propose a joint diagonalization-based algorithm (JD). To improve reconstruction error for large numbers of LoRAs while keeping the number of parameters in check, we propose a clustering approach where each cluster is compressed independently using the joint diagonalization algorithm. Our clustering algorithm is based on alternating between optimizing the cluster assignments and the per-cluster reconstruction error.
Figure 1 showcases the benefits of joint compression. When serving up to 64 unique LoRAs, we use JD without clustering and for 128 or more, we pick the number of clusters to match the performance of compressed and original LoRAs. In each case, the GPU memory footprint of the compressed and original LoRAs is matched for a fair comparison to vLLM’s multi-LoRA inference engine. When serving over 1000 LoRAs, compression increases throughput $1.6\!\times$ and maintains 80% of the throughput of serving the base LLM (or a single LoRA merged into the LLM). §6 presents detailed results.
@@
  • For large numbers of LoRAs, we scale joint compression by proposing a clustering algorithm where each cluster is jointly compressed to minimize reconstruction error.
  • We establish theoretical guarantees for the reconstruction error of our compression formulation and relate reconstruction loss to performance empirically.
  • We train a collection of more than 1000 high-quality LoRAs for Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) on 1000 natural instruction tasks (Wang et al., 2022) and demonstrate that our compression techniques preserve the performance of the original LoRAs. We will release over a 1000 LoRAs to facilitate future work as well as the code for our method.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT) has become prevalent for updating foundation models thanks to the need for efficiency in training and communication (Lialin et al., 2023). Many PEFT methods have been proposed, e.g. (Houlsby et al., 2019; Liu et al., 2022b) and LoRA (Hu et al., 2021) became the standard, partially due to the ease of switching between LoRAs in inference time.
Several works improve LoRA (Liu et al., 2024; Wang et al., 2024), sometimes with algebraic methods like SVD (Meng et al., 2024; Zhang et al., 2023; Jiang et al., 2023b) or by leveraging its statistical properties (Zhu et al., 2024; Zeng & Lee, 2024). Relatively few, however, accelerate inference times. S-LoRA (Sheng et al., 2023) provides an efficient means of switching between LoRAs. Wen & Chaudhuri (2024) adapt training to reduce batch multiplications, accelerating inference. Our method achieves a similar outcome (see Appendix D) without changing the LoRA formulation or requiring that LoRAs be trained in a dedicated way; future improvements to LoRA will also benefit from this aspect of our work (e.g., Meng et al. (2024)).
Punica (Chen et al., 2023) introduces Segmented Gather Matrix-Vector Multiplication (SGMV) to optimize multi-LoRA serving by parallelizing feature-weight multiplications in batches and grouping requests that use the same LoRA. Our approach, by contrast, reduces parameters as a means to serve multiple LoRAs efficiently, providing an orthogonal strategy that can be seamlessly integrated with Punica’s methods to enhance performance. In our vLLM experiments, we leveraged the Punica kernel for multi-LoRA implementation, demonstrating the application of our method in conjunction with Punica’s optimizations.
Other research proposes alternative PEFT methods that can be more parameter-efficient than LoRA. For example, VeRA (Kopiczko et al., 2024) fine-tunes LLMs by sharing global static parameters while learning local scaling variables; (IA)3 (Liu et al., 2022a) also reduces adapter parameter counts. However, none of these approaches has been as extensively tested or widely-adopted as LoRA. As a result, work that builds on LoRA enjoys a practical advantage due to its broad acceptance in practice.
There are many efforts to compress models (Cheng et al., 2017; Gholami et al., 2022; Sharma et al., 2024; Li et al., 2018). Predominantly, pruning and sparsification methods delete weights (Yadav et al., 2023a), and quantization methods reduce the weights’ precision (Dettmers et al., 2024). Some works compress weights to reduce model size but typically require decompression and hence do not save GPU memory (Hershcovitch et al., 2024). Similarly to our work, a few note increased performance and generalization after compression (Yadav et al., 2023a; Nadjahi et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2024).
Our work also relates to model merging (Choshen et al., 2022; Wortsman et al., 2022; Matena & Raffel, 2021) and mixtures of experts (Muqeeth et al., 2024; Yadav et al., 2024). These methods reuse models trained by others (Choshen et al., 2023; Raffel, 2023), serving them together as one compressed model. Despite this similarity, these methods create a single general model that acts on any input, while ours yields more performant per-task solutions.
@@ 3 Rank-Based LoRA Compression
While our compression technique has access only to a collection of $\{(A_{i},B_{i})\}_{i=1}^{n}$ pairs, in our experiments we will assess the efficacy of compression by comparing how the compressed matrices perform relative to uncompressed LoRAs on typical data. For this reason, although in this section we optimize a Frobenius norm reconstruction error relative to the product $B_{i}A_{i}$, this is a proxy for the nonlinear and complex way that compression errors in the adapters impact transformer performance. Our experiments will thus focus on the performance of the compressed LoRAs against the uncompressed versions on real data in §6.
Our compression methods significantly reduce the overall number of parameters. Reducing parameters theoretically accelerates storage and serving of a collection of LoRAs. This reduction, however, alters the computational dynamics during inference, so parameter reduction alone does not immediately imply faster throughput. In light of the complexities of GPU optimization, we experimentally assess throughput under realistic conditions in §6.4.
@@ 3.1 Joint Diagonalization
To scale to many LoRAs, the compressed number of parameters should not scale linearly with $n$. Hence compressing each LoRA individually (e.g., via SVD as in our experimental baselines) is inherently limited. To address this, we suggest a Joint Diagonalization (JD) method, which optimizes a shared basis onto which we can project the set of $n$ LoRAs. This allows structure to be shared, implicitly grouping and/or merging the collection of LoRAs.
In this model, each LoRA product $B_{i}A_{i}$ is factorized into the form $U\Sigma_{i}V$, where $U$ and $V$ are shared across all LoRAs and $\Sigma_{i}$ is specific to each LoRA. In this formulation, every $\Sigma_{i}$ shares the same rank $r$. This allows $U$ and $V$ to be pre-loaded onto the GPU, with $\Sigma_{i}$ loaded when necessary for each batch. The matrices $\Sigma_{i}$ can be either diagonal or small square matrices, thus significantly reducing the number of LoRA-specific parameters and accelerating multi-LoRA serving.
Objective function. Motivated by the relationship of singular value decomposition to minimizing the Frobenius norm of the reconstruction error, we also propose to minimize the Frobenius norm of the adapter matrix approximation error. Specifically, we use the following objective:
@@
Full $\Sigma_{i}$ approximation. The first method we call JD-Full. Without loss of generality, $U$ and $V$ can be constrained to be orthogonal, so long as $\Sigma_{i}$ remains an unconstrained full matrix. JD-Full adopts this restriction to make the optimization better posed, but note it does not restrict the expressiveness of the objective equation 1. This setting yields the following optimization problem:
$\operatorname{JD-Full}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})=$ $\!\underset{U^{\top}U=V^{\top}V=I_{r}}{\underset{\{\Sigma_{i}\}_{i=1}^{n}}{\operatorname{argmin}}}\!\!\!\!\!\sum_{i=1}^{n}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}$ $\qquad\qquad\qquad\qquad\qquad\qquad\ (\text{JD-Full})$ (2)
An efficient alternating algorithm to optimize this objective function can be found in Appendix A.
@@
$\operatorname{JD-Diag}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})=$ $\!\!{\underset{\{\Sigma_{i}\}_{i=1}^{n},U,V}{\operatorname{argmin}}}\sum_{i=1}^{n}\|B_{i}A_{i}-U\mathrm{diag}(\Sigma_{i})V^{\top}\|_{\mathrm{Fro}}^{2}$ $\qquad\qquad\qquad\qquad\qquad\qquad\ (\text{JD-Diag})$ (3)
Appendix A provides an efficient alternating least squares algorithm for this objective. This diagonal version has per-LoRA parameter savings when compared to JD-Full, since the diagonal $\Sigma_{i}$ only needs $r$ parameters instead of $r^{2}$.
@@ 3.2 Clustering
As the number of LoRAs $n$ grows and becomes more diverse, the rank $r$ needed for Joint Diagonalization to achieve good performance will tend to increase. This increases the size of each $\Sigma_{i}$ that needs to be stored, especially for JD-Full which will require $O(nr^{2})$ storage for these matrices. If the necessary $r$ grows proportionally to $n$, then this storage will eventually become the bottleneck.
To resolve this limitation with very large $n$, we propose to group the $n$ LoRAs into $k$ clusters $C_{j}$. Each cluster is given its own rank $r$ for JD compression, and the clusters are chosen such that the overall reconstruction error is minimized. Specifically, the overall objective is
[…]
optimized by alternating between cluster assignments and the JD of each cluster; Appendix A.3 provides details. Typically, the goal with large $n$ is to have $k$ grow with $n$ as $r$ becomes fixed. Comparing $k$ rank-$r$ JD-Full clusters to a rank-$kr$ JD-Full single cluster compression, the clustered approach requires $O(dkr+nr^{2})$ parameters, while the single-cluster approach requires $O(dkr+nk^{2}r^{2})$ parameters due to the increased sizes of the $\Sigma_{i}$s. While these two approaches have the same rank, they may have different reconstruction abilities. Empirically, we find that multiple clusters significantly aid performance for $n\geq 100$.